# SDTagNet: Leveraging Text-Annotated Navigation Maps for Online HD Map Construction

### v1

@@ A.1 Ablation studies
We conduct ablation studies to assess the contributions of individual components in the SD map encoder and the optimization strategies applied to the NLP tag embedding module. All results are reported on the Argoverse 2 dataset and are summarized in Table 3, Table 4, and Table 5 respectively.
Our architectural ablation studies of the SD map encoder in Table 3 reveal that only the full configuration of SDTagNet, combining all proposed modifications, yields significant performance gains compared to the SMERF [20] baseline. The results also highlight the importance of ORF identifiers, without which the performance even drops when switching to point level queries. This suggests that the provided element identifiers are essential for enabling effective point-level queries.
To investigate the viability of fusion approaches for the map decoder connection as proposed by TopoSD [27], we further investigated a combination of direct cross-attention to the SD map encoder tokens and cross-attention to additional SD map BEV features, shown as “+ BEV Ft.” in Table 3. The SD map BEV features are rasterized similarly to PMapNet [9]; however, in our case they contain the processed SD map encoder tokens instead of a regular grayscale image. Even though there are performance differences in individual classes, we could not observe a total performance gain with this combined connection method. Compared to [27], our results indicate that, when evaluated on a geographic split and with the more diverse elements of M3TR [8], regular cross-attention is sufficiently expressive on its own.
For the NLP tag embedding module, we observe in Table 4 that fine-tuning the pretrained language model with a small learning rate is crucial to align the self-supervised embeddings with the specific objectives of the downstream task. Additionally, removing our custom contrastive pretraining objective results in degraded performance. This confirms the utility of our training scheme, which guides the model to focus on semantically meaningful tag representations.
Additionally, we tested different numbers of attention heads and different values of dropout for the SD map encoder cross-attention in Table 5, settling on 8 attention heads and a dropout of 0.1 as the configuration with the best performance.
Base Enc.: SMERF [20] (all info.) AP
Pt. Lvl. NLP Emb. ORF Id. + BEV Ft. mAP dsh. sol. bou. cen. ped.
- - - - 45.9 38.0 54.4 50.0 48.5 38.2
$\checkmark$ - - - 44.3 34.9 53.0 49.5 45.5 38.7
- $\checkmark$ - - 45.8 37.0 55.4 47.9 49.2 39.3
$\checkmark$ - $\checkmark$ - 46.3 39.3 53.3 49.6 48.5 40.8
$\checkmark$ $\checkmark$ - - 45.0 33.6 53.2 50.2 47.7 40.6
$\checkmark$ $\checkmark$ $\checkmark$ - 48.1 36.0 55.2 53.3 52.6 43.3
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 48.1 38.2 56.3 52.6 50.7 42.3
@@
Frozen $\checkmark$ 45.6 36.5 54.2 49.8 48.4 39.2
$\times$ 0.5 $\checkmark$ 47.1 35.7 53.5 52.0 50.0 44.3
$\times$ 0.1 - 47.5 35.2 54.5 51.9 50.9 45.0
$\times$ 0.1 $\checkmark$ 48.1 36.0 55.2 53.3 52.6 43.3
@@ A.4 Argoverse 2 Original Split
To examine the differences to a non-geographic split setting, we furthermore evaluated a reduced model set of MapTRv2, SMERF (all info.), and SDTagNet on the original Argoverse 2 split with about 40% geographic overlap [17]. The results in Table 6 show that SDTagNet maintains superior performance in both range settings, with a much higher absolute mAP gain in the far range of +9.1 mAP compared to SMERF. It worth noting that the base performance is already 15 to 20 mAP higher in both settings, confirming the findings in [17] and indicating that significant geographic overfitting still takes place in the training of online HD map construction models.
@@
Dataset: Argoverse 2, og. split Near Range (60 m $\times$ 30 m)
@@ A.5 SD Map Augmentation Methods
0.5 0.4 - $\times$ 0.1 44.7 37.4 54.2 47.5 46.7 37.9
0.5 0.6 $\checkmark$ $\times$ 0.1 47.2 36.2 54.6 51.6 49.2 44.7
0.5 0.6 $\checkmark$ $\times$ 0.5 47.4 34.7 55.3 52.1 51.5 43.2
- - - $\times$ 0.1 48.1 36.0 55.2 53.3 52.6 43.3
@@
We focused on two main augmentation modes: Position noise and element dropping, shown in Table 7, and text annotation tag masking, shown in table 8. Unfortunately, in both cases, we were unable to note a performance benefit when evaluated on the validation set without any added noise. For the element-level noise modes in Table 7, both augmentation variants worsened performance instead, even when all SD map elements in one frame were moved with the same offset to simulate localization noise. The tag masking augmentation in table 8 uses an element augmentation rate to randomly select what percentage of elements will be modified and a tag drop rate to select how many tags will be dropped. An additional parameter considered whether only non-relevant tags are dropped or not. The performance of this augmentation mode is higher than with position noise, especially when only non-relevant tags are dropped. Nonetheless, tag masking was also not able to improve performance compared to the baseline without any augmentation. We believe that a method to prevent geographic overfitting via augmentation techniques or in other ways would significantly increase generalization and overall performance in online HD map construction, particularly when priors are also available.
A.6 Additional Qualitative Examples and Visualizations
@@
To provide a better understanding of the format and content of text annotations in OpenStreetMap, we include a larger resolution example of the OSM elements for a frame in fig. 7. Every element has annotated information, this includes buildings and roads with their names, oneway information or number of lanes for roads, pedestrian ways, pedestrian crossing points, etc. The road annotations on the lower left also contain some of the aforementioned TIGER geodetic database identifiers.
@@
Figure 8 shows additional examples of HD and SD map discrepancies we observed in nuScenes. These consist of inconsistent road borders that include loading and parking areas, differences in annotated crossing topology and annotation ranges that are too short and leave parts of the road unannotated. We hypothesize that these differences play a role in the reduced performance of all SD map prior methods in the nuScenes near range setting, where information from the camera images is more plentiful.
@@
In fig. 9 and fig. 10 we present additional comparisons of SDTagNet with the baselines PMapNet [9] and SMERF [20] on Argoverse 2 and nuScenes, respectively. Figure 9 contains examples that show how SDTagNet can utilize text-annotated information such as oneway roads or the number of lanes and transfer this into improved prediction results. Figure 10 displays the increased accuracy and topological consistency of SDTagNet on nuScenes compared to previous methods.

### v2

@@ A.1 Ablation studies
Additional ablation studies concerning the optimization strategies applied to the NLP tag embedding module and fine-tuning training regimes, conducted on the Argoverse 2 dataset, are summarized in Table˜5, and Table˜6.
For the NLP tag embedding module, we observe in Table˜5 that fine-tuning the pretrained language model with a small learning rate is crucial to align the self-supervised embeddings with the specific objectives of the downstream task. Additionally, removing our custom contrastive pretraining objective results in degraded performance. This confirms the utility of our training scheme, which guides the model to focus on semantically meaningful tag representations.
Additionally, we tested different numbers of attention heads and different values of dropout for the SD map encoder cross-attention in Table˜6, settling on 8 attention heads and a dropout of 0.1 as the configuration with the best performance.
@@
Frozen 45.6 36.5 54.2 49.8 48.4 39.2
0.5 47.1 35.7 53.5 52.0 50.0 44.3
0.1 - 47.5 35.2 54.5 51.9 50.9 45.0
0.1 48.1 36.0 55.2 53.3 52.6 43.3
@@ A.4 Argoverse 2 Original Split
To examine the differences to a non-geographic split setting, we furthermore evaluated a reduced model set of MapTRv2, SMERF (all info.), and SDTagNet on the original Argoverse 2 split with about 40% geographic overlap [18]. The results in Table˜7 show that SDTagNet maintains superior performance in both range settings, with a much higher absolute mAP gain in the far range of +9.1 mAP compared to SMERF. It worth noting that the base performance is already 15 to 20 mAP higher in both settings, confirming the findings in [18] and indicating that significant geographic overfitting still takes place in the training of online HD map construction models.
@@
Dataset: Argoverse 2, og. split Near Range (60 m 30 m)
@@ A.5 SD Map Augmentation Methods
0.5 0.4 - 0.1 44.7 37.4 54.2 47.5 46.7 37.9
0.5 0.6 0.1 47.2 36.2 54.6 51.6 49.2 44.7
0.5 0.6 0.5 47.4 34.7 55.3 52.1 51.5 43.2
- - - 0.1 48.1 36.0 55.2 53.3 52.6 43.3
@@
We focused on two main augmentation modes: Position noise and element dropping, shown in Table˜8, and text annotation tag masking, shown in table˜9. Unfortunately, in both cases, we were unable to note a performance benefit when evaluated on the validation set without any added noise. For the element-level noise modes in Table˜8, both augmentation variants worsened performance instead, even when all SD map elements in one frame were moved with the same offset to simulate localization noise. The tag masking augmentation in table˜9 uses an element augmentation rate to randomly select what percentage of elements will be modified and a tag drop rate to select how many tags will be dropped. An additional parameter considered whether only non-relevant tags are dropped or not. The performance of this augmentation mode is higher than with position noise, especially when only non-relevant tags are dropped. Nonetheless, tag masking was also not able to improve performance compared to the baseline without any augmentation. We believe that a method to prevent geographic overfitting via augmentation techniques or in other ways would significantly increase generalization and overall performance in online HD map construction, particularly when priors are also available.
A.6 Inference Time and Model Size
SDTagNet is designed with real-time capability in mind: the NLP tag encoder is based on a compact BERT model (embedding dim 144), and the SD map encoder is lightweight compared to the main perception backbone. To evaluate the real time capability of SDTagNet, we have investigated its inference time in comparison with no prior and with other SD map encoders. The parameter count and FPS of SDTagNet and baselines on an NVIDIA H100 GPU can be found in Table˜10. The additional inference time for SDTagNet is less than 7 ms per frame, and the memory overhead is marginal relative to the base model. We thus consider our approach suitable for real-time deployment in autonomous driving systems, with a negligible difference to other SD map encoding methods.
Table 10: Comparison of model parameter count, FPS, and VRAM with MapTRv2 and two SD map baseline methods. All experiments are on the Argoverse 2 dataset with the far-range setting, averaged over 10 samples and tested on a NVIDIA H100 GPU.
Model Param. Count FPS VRAM [MB]
MapTRv2 [16] (no prior) 42.12 M 16.0 3654
MapTRv2 + PMapNet [9] 50.48 M 13.9 3698
MapTRv2 + SMERF [21] 44.62 M 14.9 3662
MapTRv2 + SDTagNet 50.54 M 14.5 3686
A.7 Additional Qualitative Examples and Visualizations
@@
To provide a better understanding of the format and content of text annotations in OpenStreetMap, we include a larger resolution example of the OSM elements for a frame in fig.˜7. Every element has annotated information, this includes buildings and roads with their names, oneway information or number of lanes for roads, pedestrian ways, pedestrian crossing points, etc. The road annotations on the lower left also contain some of the aforementioned TIGER geodetic database identifiers.
@@
Figure˜8 shows additional examples of HD and SD map discrepancies we observed in nuScenes. These consist of inconsistent road borders that include loading and parking areas, differences in annotated crossing topology and annotation ranges that are too short and leave parts of the road unannotated. We hypothesize that these differences play a role in the reduced performance of all SD map prior methods in the nuScenes near range setting, where information from the camera images is more plentiful.
@@
In fig.˜9 and fig.˜10 we present additional comparisons of SDTagNet with the baselines PMapNet [9] and SMERF [21] on Argoverse 2 and nuScenes, respectively. Figure˜9 contains examples that show how SDTagNet can utilize text-annotated information such as oneway roads or the number of lanes and transfer this into improved prediction results. Figure˜10 displays the increased accuracy and topological consistency of SDTagNet on nuScenes compared to previous methods.